how to find m on elasticity?
P= 38,66
Q= 88,14
m=?
Answer #1

Change in quantity=(14-88)/88=-74/88=0.84

Change in price = (66-38)/38=28/38=0.737

Elasticity=chnage in quantity/change in price=-0.84/0.737=-1.14

Hence elasticity =1.14